OpenStudy (anonymous):

The particular solution of the differential equation dy/ dt = y/ 4 for which y(0) = 20 is y = 20e^−0.25t y = 19 + e^0.25t y = 20 e^0.25t y = 20e^4t

OpenStudy (anonymous):

@iGreen @Luigi0210 @hartnn @bibby

OpenStudy (xapproachesinfinity):

you can do this one by separation

OpenStudy (anonymous):

what would i get as an answer? :)

OpenStudy (xapproachesinfinity):

\[\frac{dy}{dt}=\frac{1}{4}y \Longrightarrow \frac{dy}{y}=\frac{1}{4}x\\ \int \frac{dy}{y}=1/4\int dx \Longrightarrow \ln y=\frac{x}{4}+C \]

OpenStudy (xapproachesinfinity):

take it from there...

OpenStudy (xapproachesinfinity):

solve for y using exp properties

OpenStudy (anonymous):

i think its C, y = 20 e^0.25t

OpenStudy (xapproachesinfinity):

idk you have to solve to figure out

OpenStudy (anonymous):

y = e^(c+x/4)

OpenStudy (anonymous):

that's what i got : )

OpenStudy (xapproachesinfinity):

go further use the powers rules to get y=Ce^{x/4}

OpenStudy (xapproachesinfinity):

oh by the way that should t not x

OpenStudy (xapproachesinfinity):

just a typo in the variable we have

OpenStudy (anonymous):

y = approximately 2.7182 ^ C +0.2500x

OpenStudy (anonymous):

2.7183*

OpenStudy (xapproachesinfinity):

what are you doing?

OpenStudy (xapproachesinfinity):

\[y=e^{t/4 +C}=e^{t/4}e^C=Ce^{t/4}\] for y(t=0)=20 we plug in t=0 \[20=Ce^0=C\Longrightarrow C=20\]

OpenStudy (xapproachesinfinity):

1/4=0.25 see what equation is good

OpenStudy (anonymous):

y = 20 e^0.25t

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (anonymous):

thank u :)

OpenStudy (anonymous):

YAY! I GOT A 100% thank u!!!!!!!!!!

OpenStudy (xapproachesinfinity):

YW